# On the Performance of the Neyman Allocation with Small Pilots

### v3

@@ 6.2 Additive Regularization
$\tilde{p}^{\text{add}}=\frac{\tilde{\sigma}(1)+\tilde{\sigma}(0)\nu}{(\tilde{\sigma}(1)+{\sigma}(0))(1+\nu)}$
@@
Figure 7 presents results for $\nu\in\{0.01,0.05,0.1,0.2,0.5\}$. As we increase $\nu$, the MSE decreases for the homoskedastic region but at the cost of the more heteroskedastic region.
@@ 6.3 Exponential Regularization
$\tilde{p}^{\text{exp}}=\frac{\left(\frac{\sigma(1)}{\sigma(0)}\right)^{\tau}}{1+\left(\frac{\sigma(1)}{\sigma(0)}\right)^{\tau}}\leavevmode\nobreak\ .$
Observe that when $\tau=1$, $\tilde{p}^{\text{add}}$ is FNA. For $10\leq\tau<$ 1, $\tilde{p}^{\text{exp}}$ regularizes the FNA towards the balanced allocation. In particular, $\tau=0$, $\tilde{p}^{\text{add}}=\frac{1}{2}$.
Figure 7 presents results for $\tau\in\{0.99,0.95,0.9,0.8,0.5\}$. As we decrease $\tau$, the MSE decreases for the homoskedastic region but at the cost of the more heteroskedastic region.
Figure 8: MSE under the Infeasible Neyman Allocation (INA), Balanced Allocation (BA) and Feasible Neyman Allocation (FNA) when $\varepsilon(1),\varepsilon(0)$ has the standard Normal distribution and the standardized Chi-Squared distribution.
Figure 8: MSE under the Infeasible Neyman Allocation (INA), Balanced Allocation (BA) and Feasible Neyman Allocation (FNA) when $\varepsilon(1),\varepsilon(0)$ has the standard Normal distribution and the standardized Chi-Squared distribution.
Figure 9: MSE under the Infeasible Neyman Allocation (INA), Balanced Allocation (BA) and Feasible Neyman Allocation (FNA) when $\varepsilon(1),\varepsilon(0)$ has the standard Normal distribution and the standardized Chi-Squared distribution.
@@ 6.4 Comparison of Solutions
As the above sections demonstrate, each potential solution entails different trade-offs between MSE in the homoskedastic and heteroskedastic region, and which vary depending on the tuning parameter chosen. In a crude attempt at comparing these methods, we consider the following. Choose tuning parameter values so that the MSEs of the methods are made as close as possible in the homoskedastic region. Then compare MSEs in the heteroskedastic region.
One such comparison is displayed in Figure 9. Here, we see that having fixed performance in the homoskedastic region, exponential regularization seems to do a better job at the heteroskedastic region. This seems to extend to the Pareto distributions, as well as other choices of tuning parameters that keep the solutions appropriately comparable. We therefore speculate that exponential regularization is an effective method for improving the small sample performance of the FNA.
@@ 7 Conclusion
We study the properties of the Feasible Neyman Allocation (FNA) in an asymptotic framework for two-wave experiments that takes pilot size to be fixed as the size of the main wave tends to infinity. In this setting, the estimated allocation has error that is not negligible even in the limit. Our asymptotic model therefore corresponds more closely to the finite sample statistical problem when pilots are small and the optimal allocation may be poorly estimated. We characterize conditions under which the difference-in-means estimator has larger asymptotic variance compared to balanced randomization, where half of the main wave is assigned to treatment and the remainder to control. This happens when the potential outcomes are relatively homoskedastic with respect to treatment status or exhibit high kurtosis – situations that may arise in practice. Our results suggest that the confluence of these factors – small pilots, homoskedasticity, heavy-tails – are disadvantageous to the FNA. Instead, researchers might find it useful to consider alternative procedures such as regularizing the FNA towards the balanced allocation.
@@ Acknowledgements
The authors would like to thank Federico Bugni, Ivan Canay, Joel Horowitz, Dean Karlan, Max Tabord-Meehan and Chris Udry for their helpful comments and suggestions. The authors report there are no competing interests to declare.
@@ Appendix A Proofs
$\mathbb{P}\left(\mathcal{L}_{m}\leq t\right)=\int_{0}^{1}\Phi\left(\frac{t}{s\left(p\right)}\right)\;G_{m}(\mathrm{d}p)$
@@
$\mathbb{E}\left[\mathcal{L}_{m}^{2}\right]=\mathbb{E}\left[s^{2}\left(\widetilde{p}\right)\right]=\mathbb{E}\left[\frac{\sigma^{2}(1)}{\widetilde{p}}+\frac{\sigma^{2}(0)}{1-\widetilde{p}}\right]$
@@
$\frac{\sigma^{2}(1)}{\widetilde{p}}+\frac{\sigma^{2}(0)}{1-\widetilde{p}}>\frac{\sigma^{2}(1)}{p_{*}}+\frac{\sigma^{2}(0)}{1-p_{*}}=\Sigma_{*}\leavevmode\nobreak\ .$
@@
$\mathbb{P}\left(\frac{\sigma^{2}(1)}{\widetilde{p}}+\frac{\sigma^{2}(0)}{1-\widetilde{p}}>\Sigma_{*}\right)>0\leavevmode\nobreak\ .$
@@
$\mathbb{E}\left[\frac{\sigma^{2}(1)}{\widetilde{p}}+\frac{\sigma^{2}(0)}{1-\widetilde{p}}\right]\geq\mathbb{E}\left[\frac{\sigma^{2}(1)}{1/2}+\frac{\sigma^{2}(0)}{1-1/2}\right]$
@@
$Z_{m}=\frac{\widehat{\sigma}(1)}{\sigma(1)}\bigg{/}\frac{\widehat{\sigma}(0)}{\sigma(0)}\leavevmode\nobreak\ .$
@@
$\Leftrightarrow$ $\frac{\sigma^{2}(1)}{\sigma^{2}(0)}-\mathbb{E}\left[\frac{1}{Z_{m}}+Z_{m}\right]\frac{\sigma(1)}{\sigma(0)}+1\leq 0$
@@
$\frac{\sigma(1)}{\sigma(0)}\in\left[B_{m}-\sqrt{B_{m}^{2}-1},B_{m}+\sqrt{B_{m}^{2}-1}\right]=C_{m}\leavevmode\nobreak\ ,$
@@
$B_{m}^{2}-(B_{m}^{2}-1)=1\Rightarrow\left(B_{m}-\sqrt{B_{m}^{2}-1}\right)\left(B_{m}+\sqrt{B_{m}^{2}-1}\right)=1\leavevmode\nobreak\ ,$
@@
$B_{m}-\sqrt{B_{m}^{2}-1}=\frac{1}{B_{m}+\sqrt{B_{m}^{2}-1}}=\frac{1}{x}\leavevmode\nobreak\ .$
@@
$2B_{m}>\frac{1}{\mathbb{E}\left[Z_{m}\right]}+\mathbb{E}\left[Z_{m}\right]\geq 2$
@@
$B_{m}-1=\frac{1}{2}\left[\frac{\sigma(1)}{\sigma(0)}\,\text{Bias}\left(\widetilde{p}\right)+\frac{\sigma(0)}{\sigma(1)}\,\text{Bias}\left(\frac{1}{\widetilde{p}}\right)\right]=W_{m}\leavevmode\nobreak\ .$
@@
$f^{\prime}(x)=1-\frac{1}{x^{2}}\quad,\quad$ $f^{\prime}(1)=0$
$f^{\prime\prime}(x)=\frac{2}{x^{3}}\quad,\quad$ $f^{\prime\prime}(1)=2$
@@
$m\left(B_{m}-1\right)\to\frac{V}{2}$ (A.1)
@@
$\sqrt{m}\left(\left(B_{m}+\sqrt{B_{m}^{2}-1}\right)-1\right)=\frac{m(B_{m}-1)}{\sqrt{m}}+\sqrt{m(B_{m}-1)\left(\frac{m(B_{m}-1)}{m}+2\right)}\to\sqrt{V}\leavevmode\nobreak\ .$
@@
$\mathcal{L}^{d}(F)\geq\left(\sigma(1)-\sigma(0)\right)^{2}\geq-K$
@@
$\mathcal{L}^{d}(F)\to-K\quad\mbox{ and }\quad\mathcal{L}^{r}(F)\to\frac{1}{2}\leavevmode\nobreak\ .$
@@
$X=\begin{cases}-1&\mbox{ w.p. }\frac{1}{2\kappa}\\ 0&\mbox{ w.p. }1-\frac{1}{\kappa}\\ 1&\mbox{ w.p. }\frac{1}{2\kappa}\end{cases}\leavevmode\nobreak\ .$
@@ Appendix B Additional Empirical Examples
B.1 [21, ]
[21, ] conduct an RCT to test the hypothesis that math game play in pre-school prepares poor children for formal math in primary school. Their study, conducted in Delhi, India, with the organization Pratham, involved 214 pre-schools with 1540 children, with treatment assigned at the school level. In the Math treatment arm, children were led by facilitators to play math games over the course of four months, while the control group received lessons according to Pratham’s usual curriculum. To distinguish the effect of the math games from the effect of engagement with adults, the experiment further involved a Social treatment arm, where social games were played. The outcomes of interest are Math Skills – subdivided into Symbolic Math Skills and Non-Symbolic Math Skills – as well as Social Skills, as measured by Pratham’s standardized tests. Since the authors are interested in persistence of treatment effects, they measure these outcomes at the following times after intervention: 0-3 months (Endline 1), 6-9 months (Endline 2) and 12-15 months (Endline 3). They find that the Math intervention has positive effects on Non-Symbolic Mathematical skills across all three Endlines, while Symbolic Mathematical skills only improves in Endline 1.
@@
Table 7: Individual Level Heteroskedasticity in [21, ]
@@
Suppose we are concerned about correlation across students in the same pre-school. We can redefine our unit of observation to be the school by taking averages across students in the same school. The Neyman Allocation then tells us how many schools to allocate to treatment. In this case, the standard deviation in the mean across schools is the relevant counterpart to $\sigma(1)$ and $\sigma(0)$. They are presented in Table 8. As before, we see that outcomes are relatively homoskedastic across schools, though less so than in the individual level case. Nonetheless, the ratio falls between $0.84$ and $1.55$. Once we consider schools to be the unit of treatment, however, the effective pilot size also shrinks, such that the drawbacks of the estimated Neyman Allocation may be even more pronounced. All in all, the [21, ] example supports our case of relative homoskedasticity in empirical applications.
Table 8: School Level Heteroskedasticity in [21, ]
@@
B.2 [22, ]
[22, ] study the Oregon Health Insurance Experiment, in which uninsured, low-income adults were randomly given the opportunity to apply for Medicaid. Over the course of a month in February 2008, Oregon conducted extensive public awareness campaign to encourage participation in the lottery. From a total of 89,824 sign-ups, 35,169 individuals (from 29,664 households) were selected. They, and any members of their households were then given the opportunity to apply for Medicaid. Hence, treatment occurred at the household level.
The authors used the data to study a variety of outcomes. In this section, we focus on their first set of results, which concern healthcare utilization. In particular, we revisit the outcome variables used in Tables V and VI of [22, ], which are obtained from survey data (as opposed to administrative data), and are hence publicly available. Inline with the authors’ results on the Intent-to-Treat effect, we define the treated group as those selected by the lottery. We note that the authors apply sampling weights to correct for differential response rates to the survey. We follow their weighting scheme in computing our results.
@@
Table 9: Individual Level Heteroskedasticity in [22, ]
@@
Table 10: Household Level Heteroskedasticity in [22, ]
@@
B.3 [29, ]
Business plan competitions are growing in popularity as a way of fostering high growth entrepreneurship in developing countries. [29, ] studies the Youth Enterprise With Innovation in Nigeria (YouWiN!) program, which distributed up to US\$64,000 to winners. A portion of the awards were reserved for business plans that were clearly superior to the rest. 1,841 entrepreneurs, determined to be of medium quality, were entered into a lottery, from which 729 were selected for the award. Three rounds of surveys were then conducted at 1,2 and 3 years after the application respectively.
We focus on the first set of results in [29, ] – presented in Table 2 – concerning the effect of the award on start-up and survival. Here, they find that the grant persistently increased the probability that the entrepreneur was operating a firm, as well as the number of hours they spent in self-employment. We present the standard deviations of these outcome variables in Table 11. Here we see that hours in self employment is roughly homoskedastic across all three periods. However, the outcome on whether the entrepreneur is operating a firm is arguably highly heteroskedastic, with standard deviations that is as small as 0.44 that of the control group. As in our earlier example, we find high kurtosis in these outcome variables, displayed in Table 12.
We do not estimate $C_{m}$ in this example because almost all entrepreneurs in the treated group operate their own firms in the sample. As such, a small random sub-sample (e.g. of size below 200) from this group has variance 0 with high probability, impeding the estimation of $C_{m}$. These pathological cases are revealing. In a pilot, if the treated group has variance 00 in the outcome, the estimated Neyman Allocation assigns 00 units to treatment in the full experiment. The high probability of such an “extreme” outcome with small pilots is precisely the danger which we are warning against. We conclude that the estimated Neyman from a small pilot will likely lead to adverse results given the DGP in [29, ].
Table 11: Heteroskedasticity in [29, ]

### v4

@@ 6.2 Additive Regularization
$\tilde{p}^{\text{add}}=\frac{\tilde{\sigma}(1)+\tilde{\sigma}(0)\nu}{(\tilde{\sigma}(1)+\tilde{\sigma}(0))(1+\nu)}~{}.$
@@
Figure 7 presents results for $\nu\in\{0.01,0.05,0.1,0.2,0.5\}$. As we increase $\nu$, the MSE decreases for the homoskedastic region but increases for the heteroskedastic region.
@@ 6.3 Exponential Regularization
$\tilde{p}^{\text{exp}}=\frac{\left(\frac{\tilde{\sigma}(1)}{\tilde{\sigma}(0)}\right)^{\tau}}{1+\left(\frac{\tilde{\sigma}(1)}{\tilde{\sigma}(0)}\right)^{\tau}}~{}.$
Observe that when $\tau=1$, $\tilde{p}^{\text{add}}$ is FNA. For $10\leq\tau<$ 1, $\tilde{p}^{\text{exp}}$ regularizes the FNA towards the balanced allocation. In particular, for $\tau=0$, $\tilde{p}^{\text{add}}=\frac{1}{2}$. Figure 7 presents results for $\tau\in\{0.99,0.95,0.9,0.8,0.5\}$. As before, when we decrease $\tau$, the MSE decreases for the homoskedastic region but increases for the heteroskedastic region.
Figure 8: MSE under exponential regularization. We compare the Infeasible Neyman Allocation (INA), Balanced Allocation (BA), Feasible Neyman Allocation (FNA), as well as exponential regularization with various values of $\tau$ (Exp - $\tau$).
Figure 9: MSE under various candidate solutions. The tuning parameter for each solution is chosen so that their performance in the homoskedastic region is approximately comparable.
@@ 6.4 Comparison of Solutions
As the above sections demonstrate, each potential solution entails different trade-offs between MSE in the homoskedastic and heteroskedastic region. These trade-offs also vary depending on the tuning parameter chosen. In a crude attempt at comparing these methods, we consider the following. Choose tuning parameter values so that the MSEs of the methods are made as close as possible in the homoskedastic region. Then compare MSEs in the heteroskedastic region.
One such comparison is displayed in Figure 9. Here, we see that having fixed performance in the homoskedastic region, exponential regularization seems to do a better job in the heteroskedastic region. This conclusion ostensibly extends to the Pareto distributions, as well as other choices of tuning parameters that keep the solutions appropriately comparable. We therefore speculate that exponential regularization is an effective method for improving the small sample performance of the FNA.
@@ 7 Conclusion
We study the properties of the Feasible Neyman Allocation (FNA) in an asymptotic framework for two-wave experiments that takes pilot size to be fixed as the size of the main wave tends to infinity. In this setting, the estimated allocation has error that is not negligible even in the limit. Our asymptotic model therefore corresponds more closely to the finite sample statistical problem in which pilots are small and the optimal allocation may be poorly estimated. We characterize the conditions under which the difference-in-means estimator has larger asymptotic variance compared to balanced randomization, where half of the main wave is assigned to treatment and the remainder to control. This happens when the potential outcomes are relatively homoskedastic with respect to treatment status or exhibit high kurtosis – situations that may arise in practice. Our results suggest that the confluence of these factors – small pilots, homoskedasticity, heavy-tails – are disadvantageous to the FNA. Instead, researchers might find it useful to consider alternative procedures such as regularizing the FNA towards the balanced allocation.
@@ Acknowledgements
We thank the editor, the associate editor and three anonymous referees for comments and suggestions that greatly improved the article. We are also grateful to Federico Bugni, Ivan Canay, Joel Horowitz, Dean Karlan, Max Tabord-Meehan and Chris Udry for helpful discussions. We report that there are no competing interests to declare.
@@ Appendix A Proofs
$\mathbb{P}\left(\mathcal{L}_{m}\leq t\right)=\int_{0}^{1}\Phi\left(\frac{t}{s\left(p\right)}\right)\;G_{m}(\mathrm{d}p)~{}.$
@@
$\mathbb{E}\left[\mathcal{L}_{m}^{2}\right]=\mathbb{E}\left[s^{2}\left(\widetilde{p}\right)\right]=\mathbb{E}\left[\frac{\sigma^{2}(1)}{\widetilde{p}}+\frac{\sigma^{2}(0)}{1-\widetilde{p}}\right]~{}.$
@@
$\frac{\sigma^{2}(1)}{\widetilde{p}}+\frac{\sigma^{2}(0)}{1-\widetilde{p}}>\frac{\sigma^{2}(1)}{p_{*}}+\frac{\sigma^{2}(0)}{1-p_{*}}=\Sigma_{*}~{}.$
@@
$\mathbb{P}\left(\frac{\sigma^{2}(1)}{\widetilde{p}}+\frac{\sigma^{2}(0)}{1-\widetilde{p}}>\Sigma_{*}\right)>0~{}.$
@@
$\mathbb{E}\left[\frac{\sigma^{2}(1)}{\widetilde{p}}+\frac{\sigma^{2}(0)}{1-\widetilde{p}}\right]\geq\mathbb{E}\left[\frac{\sigma^{2}(1)}{1/2}+\frac{\sigma^{2}(0)}{1-1/2}\right]~{}.$
@@
$Z_{m}=\frac{\widehat{\sigma}(1)}{\sigma(1)}\bigg{/}\frac{\widehat{\sigma}(0)}{\sigma(0)}~{}.$
@@
$\Leftrightarrow$ $\frac{\sigma^{2}(1)}{\sigma^{2}(0)}-\mathbb{E}\left[\frac{1}{Z_{m}}+Z_{m}\right]\frac{\sigma(1)}{\sigma(0)}+1\leq 0~{}.$
@@
$\frac{\sigma(1)}{\sigma(0)}\in\left[B_{m}-\sqrt{B_{m}^{2}-1},B_{m}+\sqrt{B_{m}^{2}-1}\right]=C_{m}~{},$
@@
$B_{m}^{2}-(B_{m}^{2}-1)=1\Rightarrow\left(B_{m}-\sqrt{B_{m}^{2}-1}\right)\left(B_{m}+\sqrt{B_{m}^{2}-1}\right)=1~{},$
@@
$B_{m}-\sqrt{B_{m}^{2}-1}=\frac{1}{B_{m}+\sqrt{B_{m}^{2}-1}}=\frac{1}{x}~{}.$
@@
$2B_{m}>\frac{1}{\mathbb{E}\left[Z_{m}\right]}+\mathbb{E}\left[Z_{m}\right]\geq 2~{}.$
@@
$B_{m}-1=\frac{1}{2}\left[\frac{\sigma(1)}{\sigma(0)}\,\text{Bias}\left(\widetilde{p}\right)+\frac{\sigma(0)}{\sigma(1)}\,\text{Bias}\left(\frac{1}{\widetilde{p}}\right)\right]=W_{m}~{}.$
@@
$f^{\prime}(x)=1-\frac{1}{x^{2}}\quad,\quad$ $f^{\prime}(1)=0~{},$
$f^{\prime\prime}(x)=\frac{2}{x^{3}}\quad,\quad$ $f^{\prime\prime}(1)=2~{}.$
@@
$m\left(B_{m}-1\right)\to\frac{V}{2}~{}.$ (A.1)
@@
$\sqrt{m}\left(\left(B_{m}+\sqrt{B_{m}^{2}-1}\right)-1\right)=\frac{m(B_{m}-1)}{\sqrt{m}}+\sqrt{m(B_{m}-1)\left(\frac{m(B_{m}-1)}{m}+2\right)}\to\sqrt{V}~{}.$
@@
$\mathcal{L}^{d}(F)\geq\left(\sigma(1)-\sigma(0)\right)^{2}\geq-K~{}.$
@@
$\mathcal{L}^{d}(F)\to-K\quad\mbox{ and }\quad\mathcal{L}^{r}(F)\to\frac{1}{2}~{}.$
@@
$X=\begin{cases}-1&\mbox{ w.p. }\frac{1}{2\kappa}\\ 0&\mbox{ w.p. }1-\frac{1}{\kappa}\\ 1&\mbox{ w.p. }\frac{1}{2\kappa}\end{cases}~{}.$
@@ Appendix B Additional Empirical Examples
  • Section B.1: Dillon et al. (2017)
  • Section B.2: Finkelstein et al. (2012)
  • Section B.3: McKenzie (2017)
  • Section B.4: Chong et al. (2015)
  • Section B.5: Deming et al. (2016)
  • Section B.6: Bloom et al. (2014)
  • Section B.7: Galiani and McEwan (2013)
  • Section B.8: Bryan et al. (2015)
B.1 Dillon et al. (2017)
Dillon et al. (2017) conduct an RCT to test the hypothesis that math game play in pre-school prepares poor children for formal math in primary school. Their study, conducted in Delhi, India, with the organization Pratham, involved 214 pre-schools with 1540 children, with treatment assigned at the school level. In the Math treatment arm, children were led by facilitators to play math games over the course of four months, while the control group received lessons according to Pratham’s usual curriculum. To distinguish the effect of the math games from the effect of engagement with adults, the experiment further involved a Social treatment arm, where social games were played. The outcomes of interest are Math Skills – subdivided into Symbolic Math Skills and Non-Symbolic Math Skills – as well as Social Skills, as measured by Pratham’s standardized tests. Since the authors are interested in persistence of treatment effects, they measure these outcomes at the following times after intervention: 0-3 months (Endline 1), 6-9 months (Endline 2) and 12-15 months (Endline 3). They find that the Math intervention has positive effects on Non-Symbolic Mathematical skills across all three Endlines, while Symbolic Mathematical skills only improves in Endline 1.
@@
Table 7: Individual Level Heteroskedasticity in Dillon et al. (2017)
@@
Suppose we are concerned about correlation across students in the same pre-school. We can redefine our unit of observation to be the school by taking averages across students in the same school. The Neyman Allocation then tells us how many schools to allocate to treatment. In this case, the standard deviation in the mean across schools is the relevant counterpart to $\sigma(1)$ and $\sigma(0)$. They are presented in Table 8. As before, we see that outcomes are relatively homoskedastic across schools, though less so than in the individual level case. Nonetheless, the ratio falls between $0.84$ and $1.55$. Once we consider schools to be the unit of treatment, however, the effective pilot size also shrinks, such that the drawbacks of the estimated Neyman Allocation may be even more pronounced. All in all, the Dillon et al. (2017) example supports our case of relative homoskedasticity in empirical applications.
Table 8: School Level Heteroskedasticity in Dillon et al. (2017)
@@
B.2 Finkelstein et al. (2012)
Finkelstein et al. (2012) study the Oregon Health Insurance Experiment, in which uninsured, low-income adults were randomly given the opportunity to apply for Medicaid. Over the course of a month in February 2008, Oregon conducted extensive public awareness campaign to encourage participation in the lottery. From a total of 89,824 sign-ups, 35,169 individuals (from 29,664 households) were selected. They, and any members of their households were then given the opportunity to apply for Medicaid. Hence, treatment occurred at the household level.
The authors used the data to study a variety of outcomes. In this section, we focus on their first set of results, which concern healthcare utilization. In particular, we revisit the outcome variables used in Tables V and VI of Finkelstein et al. (2012), which are obtained from survey data (as opposed to administrative data), and are hence publicly available. Inline with the authors’ results on the Intent-to-Treat effect, we define the treated group as those selected by the lottery. We note that the authors apply sampling weights to correct for differential response rates to the survey. We follow their weighting scheme in computing our results.
@@
Table 9: Individual Level Heteroskedasticity in Finkelstein et al. (2012)
@@
Table 10: Household Level Heteroskedasticity in Finkelstein et al. (2012)
@@
B.3 McKenzie (2017)
Business plan competitions are growing in popularity as a way of fostering high growth entrepreneurship in developing countries. McKenzie (2017) studies the Youth Enterprise With Innovation in Nigeria (YouWiN!) program, which distributed up to US\$64,000 to winners. A portion of the awards were reserved for business plans that were clearly superior to the rest. 1,841 entrepreneurs, determined to be of medium quality, were entered into a lottery, from which 729 were selected for the award. Three rounds of surveys were then conducted at 1,2 and 3 years after the application respectively.
We focus on the first set of results in McKenzie (2017) – presented in Table 2 – concerning the effect of the award on start-up and survival. Here, they find that the grant persistently increased the probability that the entrepreneur was operating a firm, as well as the number of hours they spent in self-employment. We present the standard deviations of these outcome variables in Table 11. Here we see that hours in self employment is roughly homoskedastic across all three periods. However, the outcome on whether the entrepreneur is operating a firm is arguably highly heteroskedastic, with standard deviations that is as small as 0.44 that of the control group. As in our earlier example, we find high kurtosis in these outcome variables, displayed in Table 12.
We do not estimate $C_{m}$ in this example because almost all entrepreneurs in the treated group operate their own firms in the sample. As such, a small random sub-sample (e.g. of size below 200) from this group has variance 0 with high probability, impeding the estimation of $C_{m}$. These pathological cases are revealing. In a pilot, if the treated group has variance 00 in the outcome, the estimated Neyman Allocation assigns 00 units to treatment in the full experiment. The high probability of such an “extreme" outcome with small pilots is precisely the danger which we are warning against. We conclude that the estimated Neyman from a small pilot will likely lead to adverse results given the DGP in McKenzie (2017).
Table 11: Heteroskedasticity in McKenzie (2017)